# Hierarchical Mixture of Experts: Generalizable Learning for High-Level Synthesis

### v3

@@ Methods
Third, we conduct an ablation study on the two-stage training and the high-level gating network’s design, and the results are listed in Table 6. There are two components in the two-stage training: the warmup epochs, and training three experts jointly and separately in turn after the warmup. We either disable the alternative training or the warmup. The lowest MSE is achieved when we use both of them. Also, the high-level gating network based on hidden representations is better than that based on input features.
More ablation studies are in the appendix. They verify that the hierarchical MoE’s performance is not due to increased parameter size or expert number, but the hierarchical structure and combination of three granularities.
@@
We want to unveil the mystery of the gating networks. We show the average assigned weights by the high-level gating network in Table 7. According to Table 1, “Fd”, “Gemv” and “Tr” have a strong preference for a certain granularity, while the other three kernels do not. Among them, the block MoE performs the best for “Fd”, and it is also assigned the highest weight; graph MoE is the best expert for “Gemv” and “Tr”, and it is also assigned the highest or nearly the highest weight. The weights are partially explainable.
@@
We pick the best hierarchical MoE model based on the previous five repeated offline evaluation experiments to do a case study on the low-level MoE. Due to the limited space, here we only analyze the block MoE, and we analyze the node and graph MoEs in the appendix. The pragmas modify the pseudo nodes, so block MoE is a window for us to analyze the pragmas. We summarize the weights of each expert for each pragma type in Table 8. There are three types of pragmas: loop tiling, pipeline, and parallelization. The third expert is good at dealing with loop tiling, pipeline, and small parallel factors; the second expert is good at dealing with large parallel factors; the other two experts deal more with pseudo nodes that do not have pragmas. Different experts diversify their roles, which improves generalizability.
@@ Conclusion
Domain generalization is a big challenge for HLS prediction models. Based on the unique challenges and opportunities, we propose the hierarchical MoE structure. In the low-level MoE, we apply MoE on one of the three natural granularities of the graph: node, basic block, or graph. In the high-level MoE, we aggregate the three low-level MoE models, so that different data points can flexibly decide which one to use. To address the severe expert polarization, we propose a two-stage training strategy. Extensive experiments have verified its effectiveness. Nonetheless, the generalizability of HLS prediction models still remains a big challenge, and future works can further improve the performance.
@@ Appendix A Additional Background Knowledge of HLS
As the HLS prediction task is quite new to the AI community, we would like to illustrate this task at the beginning of the appendix. Taking the “syr2k” kernel as an example, Figure 3 shows its source code. As explained in the Introduction section, the source code consists of the program that describes the FPGA’s function, which is similar to any normal C program, and several pragmas. Now the pragmas are represented by the placeholders “auto{pragma index}”. We utilize the machine learning model to find the best pragmas to insert in these placeholders.
@@ Appendix B Additional Problem Analysis
Nonetheless, a unique opportunity is that some different kernels share some similar substructions, where both their similarities and differences should be carefully noticed. The similarity could benefit the MoE model by utilizing the knowledge learned from other kernels. Nonetheless, the difference is also very important. An example is the “syr2k” kernel and the “syrk” kernel. They only differ in one line inside a nested loop, shown in Figure 5 and Figure 3. However, their best pragma designs are very different. In the best pragma design of “syrk”, the tiling factor of “__TILE__L2” is 60, but it is 1 for “syr2k”; the parallel factor of “__PARA__L1” and “__PARA__L2” are 20 and 1 for “syrk”, but they are 8 and 4 for “syr2k”. It is challenging for the model to not only benefit from the similarities and also observe the slight differences. Therefore, domain generalization is difficult for HLS prediction models. The MoE structure can utilize similar experts to deal with the similarities, but it can also discern the differences and use different experts to deal with them.
@@ Appendix C Detailed Experiment Settings
We also construct three new kernels by ourselves. The first kernel, “att-3mm”, calculates the attention mechanism widely used in the Transformer (Vaswani et al. 2023) architecture. Its input is an embedding matrix and two parameter matrices. It uses the embedding matrix and the first parameter matrix to calculate the Q matrix, uses the embedding matrix and the second parameter matrix to calculate the K matrix, and then multiplies Q and K to get the attention matrix. The second kernel, “att-3mm-fuse”, does the same calculation but it fuses the outer loop. The third kernel, “vmmv”, is constructed by combining vector-matrix multiplication and matrix-matrix multiplication inspired by the “atax” kernel and the “gemm-p” kernel in the HLSyn benchmark.
@@
Table 10: Dataset statistics of the complex kernels. “#Regre” denotes the number of regression data. The kernel acronyms represent “3d-rendering”, “spam-filter”, “att-3mm”, “att-3mm-fuse”, and “vmmv”.
@@ Appendix D Supplementary Experiment Results
For graph MoE, based on the expert assignment weights of each data point, we calculate the average expert assignment weights of each kernel. Then, we can calculate the cosine similarity between two kernels using the weights. Table 16 lists the most similar source kernel for each target kernel. “Gemver-medium” and its most similar kernel, “gesummv-medium”, are both matrix-vector multiplications; “syr2k” and its most similar kernel, “syrk’, are both symmetric rank operations; “trmm-opt” and its most similar kernel, “symm”, are both matrix multiplications. This indicates that the expert weights assigned by the graph MoE are partially explainable.
@@ Appendix E Other Detailed Information
Reproducibility. We have publicized our codes at https://github.com/weikai-li/HierarchicalMoE. The readme file contains instructions on how to run the codes. We have introduced the important hyper-parameters in this appendix, while other hyper-parameters can be found in the “src/config.py”, We use NVIDIA L40S (40GB) to train our model. It takes about 5 hours to train the model on the source kernels and about 1 hour to fine-tune it on the target kernels. Training costs about 30 GB of memory and fine-tuning costs about 12 GB of memory. We load the whole dataset into the memory before training. If the memory is not sufficient, we have an argument to disable this feature conveniently, and the memory cost for training can be reduced to about 15 GB. We run HLS on an AMD EPYC 7V13 64-core processor. We run HLS for the selected top-10 design points in parallel. It takes about 10 CPUs, 50 GB of memory, and 3 hours to run 10 design points for one kernel.
@@ Appendix F Limitations and Future Work
Though we have made much progress on domain generalization for HLS prediction, domain generalization is still a big challenge for HLS prediction models. First, the hierarchical MoE model does not have a huge advantage over HARP on every kernel. As shown in Table 1 in the main paper, it is significantly better than HARP on “Fd” and “Gemv”, slightly better on “Gemm” and “Tr”, but it is the same as HARP or worse than HARP on “Sy” and “Ja”. This might be due to the inherent difficulty of this problem. Future research is needed to develop even stronger models.

### v4

@@ Methods
Third, we want to verify whether the two-stage training and the high-level gating network’s second design are effective. The results are reported in Table 6. The two-stage training contains two stages: (1) the warmup stage where we train the three experts separately, and (2) training the three experts jointly and separately in turn. If we disable either one of them, the loss will increase. Also, the high-level gating network’s second design (based on hidden representations) is better than its first design (based on input features).
More ablation studies are in the appendix. They verify that the hierarchical MoE’s performance is not due to increased parameter size or expert number, but due to the hierarchical structure and combination of three granularities.
@@
We want to unveil the mystery of the gating networks. We find that their output weights are partially explainable to the model’s performance. We show the average assigned weights by the high-level gating network in Table 7. According to Table 1, “Fd,” “Gemv,” and “Tr” have a strong preference for a certain granularity, while the other three kernels do not. Among them, the block MoE performs the best for “Fd” and it is also assigned the highest weight; graph MoE is the best expert for “Gemv” and “Tr,” and it is also assigned the highest or nearly the highest weight.
@@
Besides, we find that low-level gating networks are also explainable. We pick the best hierarchical MoE model based on the previous five repeated offline evaluation experiments to do a case study on the low-level MoE. Due to the limited space, here we only analyze the block MoE, and we analyze the node and graph MoEs in the appendix. The pragmas modify the pseudo nodes, so block MoE is a window for us to analyze the pragmas. We summarize the weights of each expert for each pragma type in Table 8. There are three types of pragmas: loop tiling, pipeline, and parallelization. The third expert is good at dealing with loop tiling, pipeline, and small parallel factors; the second expert is good at dealing with large parallel factors; the other two experts deal more with pseudo nodes that do not have pragmas. Different experts diversify their roles, which improves generalizability.
@@ Conclusion
Domain generalization is a big challenge for HLS prediction models. We propose the hierarchical MoE structure. In the low-level MoE, we apply MoE on one of the three natural granularities of the graph: node, basic block, or graph. In the high-level MoE, we aggregate the three low-level MoE models, so that different data points can flexibly use one of them. To address the expert polarization, we propose a two-stage training strategy. Extensive experiments have verified our model’s effectiveness. Meanwhile, the generalizability of HLS prediction models still remains a big challenge, and future works should further improve the performance.
@@ Appendix A Additional Background Knowledge of HLS
As the HLS prediction task is quite new to the AI community, we would like to illustrate this task at the beginning of the appendix. Taking the “syr2k” kernel as an example, Figure 3 shows its source code. As explained in the Introduction section, the source code consists of the program that describes the FPGA’s function, which is similar to any normal C program, and several pragmas. Now the pragmas are represented by the placeholders “auto{pragma index}.” We utilize the machine learning model to find the best pragmas to insert in these placeholders.
@@ Appendix B Additional Problem Analysis
Nonetheless, a unique opportunity is that some different kernels share some similar substructions, where both their similarities and differences should be carefully noticed. The similarity could benefit the MoE model by utilizing the knowledge learned from other kernels. Nonetheless, the difference is also very important. An example is the “syr2k” kernel and the “syrk” kernel. They only differ in one line inside a nested loop, shown in Figure 5 and Figure 3. However, their best pragma designs are very different. In the best pragma design of “syrk,” the tiling factor of “__TILE__L2” is 60, but it is 1 for “syr2k”; the parallel factor of “__PARA__L1” and “__PARA__L2” are 20 and 1 for “syrk,” but they are 8 and 4 for “syr2k.” It is challenging for the model to not only benefit from the similarities and also observe the slight differences. Therefore, domain generalization is difficult for HLS prediction models. The MoE structure can utilize similar experts to deal with the similarities, but it can also discern the differences and use different experts to deal with them.
@@ Appendix C Detailed Experiment Settings
We also construct three new kernels by ourselves. The first kernel, “att-3mm,” calculates the attention mechanism widely used in the Transformer (Vaswani et al. 2023) architecture. Its input is an embedding matrix and two parameter matrices. It uses the embedding matrix and the first parameter matrix to calculate the Q matrix, uses the embedding matrix and the second parameter matrix to calculate the K matrix, and then multiplies Q and K to get the attention matrix. The second kernel, “att-3mm-fuse,” does the same calculation but it fuses the outer loop. The third kernel, “vmmv,” is constructed by combining vector-matrix multiplication and matrix-matrix multiplication inspired by the “atax” kernel and the “gemm-p” kernel in the HLSyn benchmark.
@@
Table 10: Dataset statistics of the complex kernels. “#Regre” denotes the number of regression data. The kernel acronyms represent “3d-rendering,” “spam-filter,” “att-3mm,” “att-3mm-fuse,” and “vmmv.”
@@ Appendix D Supplementary Experiment Results
For graph MoE, based on the expert assignment weights of each data point, we calculate the average expert assignment weights of each kernel. Then, we can calculate the cosine similarity between two kernels using the weights. Table 16 lists the most similar source kernel for each target kernel. “Gemver-medium” and its most similar kernel, “gesummv-medium,” are both matrix-vector multiplications; “syr2k” and its most similar kernel, “syrk,” are both symmetric rank operations; “trmm-opt” and its most similar kernel, “symm,” are both matrix multiplications. This indicates that the expert weights assigned by the graph MoE are partially explainable.
@@ Appendix E Other Detailed Information
Reproducibility. We have publicized our codes at https://github.com/weikai-li/HierarchicalMoE. The readme file contains instructions on how to run the codes. We have introduced the important hyper-parameters in this appendix, while other hyper-parameters can be found in the “src/config.py,” We use NVIDIA L40S (40GB) to train our model. It takes about 5 hours to train the model on the source kernels and about 1 hour to fine-tune it on the target kernels. Training costs about 30 GB of memory and fine-tuning costs about 12 GB of memory. We load the whole dataset into the memory before training. If the memory is not sufficient, we have an argument to disable this feature conveniently, and the memory cost for training can be reduced to about 15 GB. We run HLS on an AMD EPYC 7V13 64-core processor. We run HLS for the selected top-10 design points in parallel. It takes about 10 CPUs, 50 GB of memory, and 3 hours to run 10 design points for one kernel.
@@ Appendix F Limitations and Future Work
Though we have made much progress on domain generalization for HLS prediction, domain generalization is still a big challenge for HLS prediction models. First, the hierarchical MoE model does not have a huge advantage over HARP on every kernel. As shown in Table 1 in the main paper, it is significantly better than HARP on “Fd” and “Gemv,” slightly better on “Gemm” and “Tr,” but it is the same as HARP or worse than HARP on “Sy” and “Ja.” This might be due to the inherent difficulty of this problem. Future research is needed to develop even stronger models.